# SynBrain: Enhancing Visual-to-fMRI Synthesis via Probabilistic Representation Learning

### v2

@@ Appendix E Comparison between Diffusion Transformer and S2N Mapper
To address this distribution gap, we propose replacing the diffusion prior with a one-step Semantic-to-Neural (S2N) Transformer Mapper. The resulting model, SynBrain (BrainVAE-S2N), directly maps to the fMRI latent distribution from CLIP embeddings without iterative sampling. This approach ensures distributional consistency and semantic grounding throughout the generation process.
@@
Appendix F Semantically Consistent Synthesis under Stochastic Sampling
@@
Appendix G Additional Results
G.1 Subject-specific Visual-to-fMRI Synthesis
@@
G.2 Few-shot Subject-adaptive Visual-to-fMRI Synthesis
To evaluate SynBrain’s generalization ability under few-shot settings, we perform subject-adaptive visual-to-fMRI synthesis by transferring models trained on one source subject to novel target subjects with only 1 hour of adaptation data. While the main paper reports transfer from Sub1 to Sub2, Sub5, and Sub7, here we further investigate the reverse direction: transferring from Sub2, Sub5, and Sub7 to Sub1. The goal is to assess whether SynBrain can consistently adapt to a novel subject given different source pretraining.
As shown in Table 8, all three source-subject models (Sub2→Sub1, Sub5→Sub1, Sub7→Sub1) achieve strong semantic-level decoding scores (Incep: 88.0–88.9%, CLIP: 86.7–86.9%) and comparable voxel-level alignment after limited adaptation. The fMRI-to-image retrieval accuracy based on the synthesized signal (Syn: 63.5–69.4%) also matches the performance observed in the Sub1→Others setting reported in the main text. These results confirm that SynBrain enables bidirectional subject transfer with minimal data, highlighting the robustness and adaptability of SynBrain in cross-subject low-resource adaptation scenarios.
G.3 Data Augmentation for Few-shot fMRI-to-Image Decoding
@@
G.4 Brain Functional Analysis

### v3

@@ Appendix D Comparison between Diffusion Transformer and S2N Mapper
To address this distribution gap, we propose replacing the diffusion transformer with a one-step Semantic-to-Neural (S2N) Transformer Mapper. The resulting model, SynBrain (BrainVAE-S2N), directly maps to the fMRI latent distribution from CLIP embeddings without iterative sampling. This approach ensures distributional consistency and semantic grounding throughout the generation process.
@@
Appendix E Semantically Consistent Synthesis under Stochastic Sampling
@@
Appendix F Vision Model Comparison
In this section, we evaluate the impact of visual encoder choice by replacing CLIP in SynBrain with two advanced visual foundation models, DINOv2 oquab2023dinov2 and MAE he2022masked . As shown in the table below, CLIP achieves the best overall performance across voxel-level and semantic levels. We attribute this to its vision-language contrastive objective, which provides stronger semantic supervision compared to purely visual models like DINOv2 and MAE. Nonetheless, the relatively strong performance of these alternatives indicates that SynBrain remains robust across different encoder choices.
We also include a comparison with MindSimulator (5-Trial), which likewise uses CLIP as its visual encoder. Notably, MindSimulator averages results over five generated fMRI trials to improve performance, whereas SynBrain achieves superior results with just a single trial. This suggests that SynBrain benefits not only from a strong encoder but also from a more effective model design and training strategy.
Method Visual Voxel-Level Semantic-Level (via decoding) Image Retrieval
Encoder MSE $\downarrow$ Pearson $\uparrow$ Cosine $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$ Raw $\uparrow$ Syn $\uparrow$
MindSimulator CLIP .417.417 .326.326 - $92.8\%$ $89.8\%$ .714.714 .402.402 - -
SynBrain CLIP* .079 .715 .769 $\textbf{96.7}\%$ $\textbf{95.9}\%$ .619 .350 $\textbf{94.7}\%$ $\textbf{99.3}\%$
MAE .097.097 .676.676 .737.737 $91.0\%$ $90.5\%$ .705.705 .389.389 $68.6\%$ $83.7\%$
DINOv2 .089.089 .679.679 .728.728 $92.6\%$ $91.8\%$ .701.701 .375.375 $76.3\%$ $87.6\%$
Table 7: Visual-to-fMRI synthesis performance with various visual encoders.
Appendix G Additional Baselines
Subject-specific Baselines.
Here we conduct additional comparisons with two representative baselines: (i) GNet st2023brain , and (ii) Linear Regression (LinearReg), a classical deterministic voxel-wise visual encoding model, using a simple two-layer linear architecture with a 4,096-dimensional hidden space. As shown in Table 8, SynBrain achieves consistently superior results across multiple levels, highlighting its effectiveness in fMRI synthesis with both voxel-level structure and semantic consistency.
Subject-adaptive baselines.
Taking MindSimulator bao2025mindsimulator as a cross-subject adaptation baseline is limited by two practical constraints. i) MindSimulator is not open-sourced and incorporates several hand-crafted components (e.g., resting-state initialization, custom noise injection), making it difficult to reproduce; ii) MindSimulator’s Autoencoder is subject-specific and tightly coupled to fixed voxel dimensions, rendering it incompatible with cross-subject scenarios where ROI sizes vary. These limitations prevent a direct and meaningful comparison between SynBrain and MindSimulator in the context of cross-subject adaptation.
To establish a transparent and reliable baseline, we implemented Linear Regression (LinearReg), the most classical encoding model, using a simple two-layer linear architecture. The first layer maps CLIP features to a 4,096-dimensional hidden space, and the second layer projects to subject-specific voxel responses. During few-shot adaptation, we freeze the first layer and train only the second layer using 1-hour data from the target subject. As shown in Table 1 and Table 9, SynBrain consistently outperforms Linear Regression across voxel-wise correlation, semantic alignment, and image retrieval accuracy, demonstrating strong generalization and cross-subject transferability under data-limited scenarios.
Method Voxel-Level Semantic-Level (via decoding) Image Retrieval
LinearReg .102.102 .676.676 .693.693 $85.8\%$ $83.7\%$ .759.759 .454.454 - -
GNet .092.092 .707.707 .740.740 $87.7\%$ $85.0\%$ .730.730 .428.428 - -
SynBrain .079 .715 .769 $\textbf{96.7}\%$ $\textbf{95.9}\%$ .619 .350 $\textbf{94.7}\%$ $\textbf{99.3}\%$
Table 8: Subject-specific visual-to-fMRI synthesis baseline performance on Subject 1.
Method (LinearReg) Voxel-Level Semantic-Level (via decoding) Image Retrieval
[…]
Table 9: Subject-adaptive visual-to-fMRI synthesis baseline performance using LinearReg.
Appendix H Additional Results
H.1 Subject-specific Visual-to-fMRI Synthesis
@@
H.2 Few-shot Subject-adaptive Visual-to-fMRI Synthesis
To comprehensively assess the generalization ability of SynBrain under few-shot settings, we conduct subject-adaptive visual-to-fMRI synthesis by transferring models trained on one source subject to novel target subjects using only one hour of adaptation data. While the main paper reports transfers from Sub1 to Sub2, Sub5, and Sub7, here we extend the evaluation in two directions: (i) expanding to additional subjects and (ii) performing reverse-direction transfers.
(i) Evaluation on Additional Subjects. To examine SynBrain’s ability to generalize across a more diverse population, we expanded our experiments to include all eight real subjects from the NSD dataset. The main experiments involved four subjects (Sub1, Sub2, Sub5, Sub7), whereas the remaining four (Sub3, Sub4, Sub6, Sub8), who completed fewer sessions, were incorporated for supplementary evaluation. As shown in Table 11 (Top), SynBrain maintains strong performance across these additional unseen individuals with only one hour of adaptation data, demonstrating robust generalization despite substantial inter-subject neural variability.
(ii) Reverse-Direction Subject Adaptation. We further investigated reverse transfer by adapting models pretrained on Sub2, Sub5, and Sub7 to Sub1. As presented in Table 11 (Bottom), all three source-subject models achieve consistently high semantic-level decoding accuracy (Inception: 88.0–88.9%, CLIP: 86.7–86.9%) and comparable voxel-level alignment after limited adaptation. Moreover, the fMRI-to-image retrieval accuracy based on the synthesized signals (Syn: 63.5–69.4%) closely matches the results observed in the Sub1→Others setting reported in the main text. These results confirm that SynBrain enables bidirectional cross-subject adaptation with minimal data, underscoring its robustness and adaptability in low-resource, cross-subject generalization scenarios.
H.3 Data Augmentation for Few-shot fMRI-to-Image Decoding
@@
Trade-off between Naturalistic and Synthetic Data.
Here we conduct additional experiments with extended synthetic data durations (up to 32 hours) and observed that the optimal performance is achieved with just 1 hour of data augmentation. As shown in Table 13, adding more synthetic data beyond this point leads to diminishing or even negative returns. This suggests a trade-off between real and generated data: while a small amount of high-quality synthetic data is beneficial, excessive augmentation may introduce distributional mismatch or redundancy, limiting further gains.
Method Low-Level High-Level Retrieval
MindEye2(1h) .235 .428 88.0% 93.3% 83.6% 80.8% .798 .459 94.0% 77.6%
MindEye2(1h)+DA(8h) .222 .417 88.5% 93.6% 84.9% 82.5% .795 .450 67.1% 68.8%
MindEye2(1h)+DA(16h) .229 .416 88.4% 93.4% 83.1% 81.8% .812 .457 64.8% 63.7%
MindEye2(1h)+DA(32h) .222 .410 88.4% 93.1% 82.8% 81.8% .813 .457 61.6% 61.9%
Table 13: Few-shot fMRI-to-image decoding with various hours of synthetic data.
H.4 Brain Functional Analysis